# Strict comparison in reduced group $C^{*}$-algebras

### v2

@@ 3.3. Proving strict comparison for $C^{*}_{r}(G)$
We follow the proof of [44, Theorem 1.6]. Set $Y=X\cup\{a\}$ to be our finite generating set of $\Gamma:=G*{\mathbb{Z}}$. Let $\varepsilon>0$ and let $S\subset\Gamma$ be a finite subset. Then there is some $R>0$ such that $S\subset B_{Y}(R)$. Let $S_{Y}(R)\subset{\mathbb{C}}\Gamma$ denote the elements supported on $B_{Y}(R)$ with $\ell^{1}$ norm equal to 1. Using compactness of $S_{Y}(R)$ with respect to the $\ell^{1}$ norm, we find a finite $\frac{\varepsilon}{3}$-net $\{a_{i}\}_{i\in I}$ for $S_{Y}(R)$. It now suffices to prove the existence of a homomorphism $\phi:\Gamma\to G$ such that $\phi|_{G}=\operatorname{id}_{G}$ and that, for all $i\in I,$
@@
Since $G$ is selfless, we get a function $f:{\mathbb{N}}\to{\mathbb{R}}$ such that $\liminf_{n}f(n)^{\frac{1}{n}}=1$ and homomorphisms $\phi_{n}:\Gamma\to G$ which are onto, are identity on $G,$ are injective on $B_{Y}(n),$ and such that $\phi_{n}(B_{Y}(n))\subset B_{X}(f(n))$. Using rapid decay of $G$ we get a polynomial $P$ such that $\left\|\lambda_{G}(a)\right\|\leq P(r)\|a\|_{2}$ for all $a$ supported on $B_{X}(r)$. Now, as $\liminf_{n}f(n)^{\frac{1}{n}}=1$, we may pick $m$ such that $P(f(2mR))^{\frac{1}{2m}}<1+\frac{\varepsilon}{3}$. Define $b_{i}=\phi_{2mR}(a_{i})$ for each $i\in I$.
Note that $\|\lambda_{\Gamma}(a_{i})\|^{2m}=\|\lambda_{\Gamma}(a_{i}^{*}a_{i})^{m}\|$ and similarly $\|\lambda_{G}(b_{i})\|^{2m}=\|\lambda_{G}(b_{i}^{*}b_{i})^{m}\|$. Since each $a_{i}$ is supported on $B_{Y}(R)$, each $(b_{i}^{*}b_{i})^{m}$ is supported on $B_{X}(f(2mR))$. Since $\phi_{2mR}$ is injective on $B_{Y}(2mR)$, we have that $\|(b_{i}^{*}b_{i})^{m}\|_{2}=\|(a_{i}^{*}a_{i})^{m}\|_{2}$. Therefore
@@
In the above proof, the injectivity of $\phi_{n}$ on the radius $n$ ball $B_{X\cup\{a\}}(n)$ is used only to obtain the equality $\|(b_{i}^{*}b_{i})^{m}\|_{2}=\|(a_{i}^{*}a_{i})^{m}\|_{2}$. If instead $\phi_{n}$ is $g(n)$-to-1 on the radius $n$ ball for some function $g:{\mathbb{N}}\to{\mathbb{R}}$ such that $\lim g(n)^{\frac{1}{n}}=1,$ then by the Cauchy-Schwarz inequality we would still have $\|(b_{i}^{*}b_{i})^{m}\|_{2}\leq\sqrt{g(2mR)}\|(a_{i}^{*}a_{i})^{m}\|_{2}$ and the rest of the proof would proceed as above. To summarize, in the definition of selfless, we can assert the weaker condition that $\phi_{n}$ is $g(n)$-to-1 for some subexponential function $g$. Note that in this setting, the group homomorphism $\phi=(\phi_{n})_{\mathcal{U}}:G*{\mathbb{Z}}\to G^{\mathcal{U}}$ is no longer necessarily injective, which isn’t a requirement anyway to prove Proposition 3.7. We suspect that this approach might also help resolve the question of strict comparison in broader generality.
@@
Since the map $\phi=(\phi_{n})_{\mathcal{U}}$ is a homomorphism, we immediately get the previous inequality for all $z\in{\mathbb{C}}\Gamma$. Then the inequality implies that we can extend $\phi$ continuously to all of $C_{r}^{*}(\Gamma)$; thus, $\phi$ is the desired embedding. ∎
@@
By [56, Proposition 5], it suffices to show that, for some free ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$ there is an injective $\ast$-homomorphism $\phi$ from the reduced free product $C^{*}_{r}(G)*C_{r}^{*}({\mathbb{Z}})$ to $C^{*}_{r}(G)^{\mathcal{U}}$ such that, composing with the canonical inclusion $C_{r}^{*}(G)\subset C_{r}^{*}(G)*C_{r}^{*}({\mathbb{Z}})$, we obtain the diagonal embedding $C_{r}^{*}(G)\subset C_{r}^{*}(G)^{\mathcal{U}}$. In other words, we require an existential embedding of $C^{*}_{r}(G)*C_{r}^{*}({\mathbb{Z}})$ in $C^{*}_{r}(G)^{\mathcal{U}}$. But this is precisely the content of Proposition 3.7. ∎
@@ Appendix A Quantitative version of Wenyuan Yang’s admissible path lemma
The aim of this section is to prove a quantitative version of W. Yang’s admissible path lemma [68, Corollary 3.4], which applies to our setting. For our purposes, we will take a fixed action on a hyperbolic space but vary the loxodromic element and thus the quasi-geodesic constants $\lambda$. So throughout, we track dependence on $\lambda$. The following quantitative version of the Morse lemma is present in [29, Section 5.2]; see [62, Theorem 1.1]) for an explicit polynomial.
@@
$[g]=\inf_{x\in X}d(gx,x).$
This infimum is in fact achieved at some point $x_{0}\in X$. We let $L_{g}=\cup_{n\in{\mathbb{Z}}}[g^{n}x_{0},g^{n+1}x_{0}]$. Let $\gamma=\cup_{n\in{\mathbb{Z}}}[g^{n}o,g^{n+1}o]$ be the axis of $g$ going through $o.$ Let $\lambda=\frac{d(go,o)}{\tau(g)}$ so that $\gamma$ is a $(\lambda,0)$-quasi-axis; i.e., if $m<n$ are integers then
$\ell(\gamma)=\sum_{i=m}^{n-1}d(g^{i}o,g^{i+1}o)\leq\lambda d(\gamma^{-},\gamma^{+})=\lambda d(g^{n-m}o,o).$
Note that clearly $[g]\leq d(go,o)=\tau(g)\lambda.$ Let $E(g)$ denote the maximal elementary (i.e., virtually cyclic) subgroup of $G$ containing $g.$ For sets $Y,Z\subset X$, let $\pi_{Y}(Z)$ denote the projection of $Z$ onto $Y.$ More precisely, $y\in\pi_{Y}(Z)$ if and only if $y\in Y$ and there exists $z\in Z$ such that $d(y,z)=\min\{d(y^{\prime},z):y^{\prime}\in Y\}$. For a set $Y\subset X$ let $\|Y\|$ denote its diameter; i.e., the smallest real number such that $d(y,y^{\prime})\leq\|Y\|$ for all $y,y^{\prime}\in Y.$
The following is due to Delzant. We note that the constant $D_{0}$ below depends only on the action $G\curvearrowright X$.
@@
([21, Proposition 5]) Let $G\curvearrowright X$ be an acylindrical action on a hyperbolic space and let $g\in G$ be a loxodromic element (i.e. a hyperbolic isometry). Then for all $h\not\in E(g)$,
@@
The following lemma is the quantitative version of [64, Lemma 2.23].
@@
There is a quadratic $Q_{2}$ such that whenever a loxodromic element $g$ and a point $x\in X$ gives a $(\lambda,0)$-quasi-axis $\gamma=\cup_{n\in{\mathbb{Z}}}[g^{n}x,g^{n+1}x]$ and $h\not\in E(g),$ then
@@
The following theorem, the main theorem of this section, is essentially a reproduction of [68, Corollary 3.4] but where we have tracked all of the constants and dependencies to verify they grow polynomially.
@@
Set $\lambda=\frac{d(go,o)}{\tau(g)}$. Let ${\mathcal{L}}$ be the collection of $(\lambda,0)$-quasi-geodesics in $X$; in particular, ${\mathcal{L}}$ includes all geodesics since (without loss of generality) $\lambda\geq 1$. Let ${\mathbb{X}}$ be defined by ${\mathbb{X}}=\{\cup_{n=k}^{\ell}[hg^{n},0,hg^{n+1}o]:h\not\in E(g)\}$ and note that ${\mathbb{X}}\subset{\mathcal{L}}.$ By Lemma A.1 and [9, Section 3], we have that ${\mathbb{X}}$ is $(\mu,\varepsilon)$-contracting with respect to ${\mathcal{L}}$ where $\mu,\varepsilon$ are uniformly bounded above by a quadratic $Q_{3}(\lambda)$. That is, whenever $q\in{\mathcal{L}}$ and $p\in{\mathbb{X}}$ such that $d(q,p)\geq Q_{3}(\lambda)$ then $\|\pi_{p}(q)\|<Q_{3}(\lambda)$. (In the notation of [68], $\mu(\lambda^{\prime},c^{\prime}),\varepsilon(\lambda^{\prime},c^{\prime})\leq Q_{3}(\lambda)$ for all $\lambda^{\prime},c^{\prime}\in{\mathbb{R}}$.)
We now obtain that for each $p\in{\mathbb{X}}$, $p$ is $\sigma$-quasiconvex where $\sigma(U)=\frac{3}{2}Q_{3}(\lambda)+2U$. Indeed, let $q$ be a geodesic in $X$ such that its endpoints $q^{+}$ and $q^{-}$ are in the $U$-neighborhood $N_{U}(p)$ of $p.$ Since $p$ is $(\mu,\varepsilon)$-contracting, either $d(q,p)<\mu$ (and we are done) or $\|\pi_{p}(q)\|<\varepsilon$. In particular, we have $d(q^{+},q^{-})\leq\varepsilon+2U$. Since $q$ is a geodesic, for any $x\in q$ we have $d(q^{+},x)+d(q^{-},x)=d(q^{+},q^{-})$. Therefore without loss of generality $d(q^{+},x)\leq\varepsilon/2+U$, implying that $d(x,\pi_{p}(x))<2U+\frac{3}{2}\varepsilon$. Thus $q\subset N_{\sigma(U)}(p)$.
By Lemma A.3, if $p,p^{\prime}\in{\mathbb{X}}$ are quasi-geodesics corresponding to $h,h^{\prime}$ in different left cosets of $E(g)$, then $p,p^{\prime}$ have bounded projection $T$ where $T=Q_{2}(\lambda)(1+[g])=O(\lambda^{3})$ since $[g]\leq d(go,o)=\tau(g)\lambda$ (and the stable length $\tau(g)$ is fixed). We note that in [68], the symbol $\tau$ is used for bounded projection but we have already used $\tau$ for stable length. By [68, Lemma 2.7], $p,p^{\prime}$ also have $\nu$-bounded intersection where $\nu(U)=2U+O(\lambda^{3})$.
@@
$P$ $=[o,h_{1}o]\cup[h_{1}o,h_{1}go]\cup\ldots\cup[h_{1}g^{n_{1}-1}o,h_{1}g^{n_{1}}o]\cup[h_{1}g^{n_{1}}o,h_{1},g^{n_{1}}h_{2}o]\cup\ldots$
$\ldots\cup[h_{1}g^{n_{1}}\cdots h_{k}g^{n_{k}-1}o,h_{1}g^{n_{1}}\cdots h_{m}g^{n_{m}}o],$
and set
@@
We are now ready to analyze the constants in the proof of [68, Proposition 3.3]. We have $c=0$ everywhere.
@@
$B_{\lambda,0}$ $=2\varepsilon(1,0)+2\mu(1,0)+\nu(\mu(1,0)+\sigma(0))+|g|_{S}$
@@
$\quad\quad\quad\ \ 2T+3\varepsilon(1,0)+6\mu(1,0),$
@@
Now, if we choose all of the $n_{i}$ so that $|n_{i}|$ are larger than $\lambda D=O(|g|_{S}^{6}),$ we can apply [68, Corollary 3.4] to obtain that our path $P$ is a $(\Lambda,0)$-quasi-geodesic. In particular, $P$ is nontrivial and so it must be that $h_{1}g^{n_{1}}h_{2}g^{n_{2}}\cdots h_{m}g^{n_{m}}\neq e$. (Note that in the case $n_{m}=0$, we can still apply [68, Corollary 3.4].) This finishes the proof. ∎

### v4

@@ 3.3. Proving strict comparison for $C^{*}_{r}(G)$
We follow the proof of [46, Theorem 1.6]. Set $Y=X\cup\{a\}$ to be our finite generating set of $\Gamma:=G*{\mathbb{Z}}$. Let $\varepsilon>0$ and let $S\subset\Gamma$ be a finite subset. Then there is some $R>0$ such that $S\subset B_{Y}(R)$. Let $S_{Y}(R)\subset{\mathbb{C}}\Gamma$ denote the elements supported on $B_{Y}(R)$ with $\ell^{1}$ norm equal to 1. Using compactness of $S_{Y}(R)$ with respect to the $\ell^{1}$ norm, we find a finite $\frac{\varepsilon}{3}$-net $\{a_{i}\}_{i\in I}$ for $S_{Y}(R)$. That is, every element in $S_{Y}(R)$ would be within $\varepsilon/3$ distance of one of the $a_{i}$’s. It now suffices to prove the existence of a homomorphism $\phi:\Gamma\to G$ such that $\phi|_{G}=\operatorname{id}_{G}$ and that, for all $i\in I,$
@@
Since $G$ is selfless, we get a function $f:{\mathbb{N}}\to{\mathbb{R}}_{+}$ such that $\liminf_{n}f(n)^{\frac{1}{n}}=1$ and homomorphisms $\phi_{n}:\Gamma\to G$ which are epimorphisms, are identity on $G,$ are injective on $B_{Y}(n),$ and such that $\phi_{n}(B_{Y}(n))\subset B_{X}(f(n))$. Using rapid decay of $G$ we get a polynomial $P$ such that $\left\|\lambda_{G}(a)\right\|\leq P(r)\|a\|_{2}$ for all $a$ supported on $B_{X}(r)$. Now, as $\liminf_{n}f(n)^{\frac{1}{n}}=1$, we may pick $m$ such that $P(f(2mR))^{\frac{1}{2m}}<1+\frac{\varepsilon}{3}$. Define $b_{i}=\phi_{2mR}(a_{i})$ for each $i\in I$.
Note that $\|\lambda_{\Gamma}(a_{i})\|^{2m}=\|\lambda_{\Gamma}(a_{i}^{*}a_{i})^{m}\|$ and similarly $\|\lambda_{G}(b_{i})\|^{2m}=\|\lambda_{G}(b_{i}^{*}b_{i})^{m}\|$. Since each $a_{i}$ is supported on $B_{Y}(R)$, $(a_{i}^{*}a_{i})^{m}$ is supported on $B_{Y}(2mR)$, and hence each $(b_{i}^{*}b_{i})^{m}=\phi_{2mR}((a_{i}^{*}a_{i})^{m})$ is supported on $B_{X}(f(2mR))$. Since $\phi_{2mR}$ is injective on $B_{Y}(2mR)$, we have that $\|(b_{i}^{*}b_{i})^{m}\|_{2}=\|(a_{i}^{*}a_{i})^{m}\|_{2}$. Therefore
@@
In the above proof, the injectivity of $\phi_{n}$ on the radius $n$ ball $B_{X\cup\{a\}}(n)$ is used only to obtain the equality $\|(b_{i}^{*}b_{i})^{m}\|_{2}=\|(a_{i}^{*}a_{i})^{m}\|_{2}$. If instead $\phi_{n}$ is $g(n)$-to-1 on the radius $n$ ball for some function $g:{\mathbb{N}}\to{\mathbb{N}}$ such that $\lim g(n)^{\frac{1}{n}}=1,$ then by the Cauchy-Schwarz inequality we would still have $\|(b_{i}^{*}b_{i})^{m}\|_{2}\leq\sqrt{g(2mR)}\|(a_{i}^{*}a_{i})^{m}\|_{2}$ and the rest of the proof would proceed as above. To summarize, in the definition of selfless, we can assert the weaker condition that $\phi_{n}$ is $g(n)$-to-1 for some subexponential function $g$. Note that in this setting, the group homomorphism $\phi=(\phi_{n})_{\mathcal{U}}:G*{\mathbb{Z}}\to G^{\mathcal{U}}$ is no longer necessarily injective, which isn’t a requirement anyway to prove Proposition 3.7. We suspect that this approach might also help resolve the question of strict comparison in broader generality.
@@
Since the map $\phi=(\phi_{n})_{\mathcal{U}}$ is a homomorphism, we immediately get the previous inequality for all $z\in{\mathbb{C}}\Gamma$. Then the inequality implies that we can extend $\phi$ continuously to all of $C_{r}^{*}(\Gamma)$; thus, $\phi$ is the desired *-homomorphism. ∎
@@
By [59, Theorem 2.6], it suffices to show that, for some free ultrafilter $\mathcal{U}$} on $\mathbb{N}$} there is an injective $\ast$-homomorphism $\phi$ from the reduced free product $C^{*}_{r}(G)*C_{r}^{*}({\mathbb{Z}})$ to $C^{*}_{r}(G)^{\mathcal{U}}$ such that, composing with the canonical inclusion $C_{r}^{*}(G)\subset C_{r}^{*}(G)*C_{r}^{*}({\mathbb{Z}})$, we obtain the diagonal embedding $C_{r}^{*}(G)\subset C_{r}^{*}(G)^{\mathcal{U}}$. The content of Proposition 3.7 is precisely producing such a *-homomorphism, and it is injective because $G*{\mathbb{Z}}$ is $C^{*}$-simple for any infinite group $G$, (see for example Corollary 3 in [23]), and as a consequence, any *-homomorphism with domain $C^{*}_{r}(G*{\mathbb{Z}})$ will be automatically injective. ∎
@@
Proof of Theorem B.
If $G$ is as in the hypothesis of Theorem B, then by Theorem 3.3 and Theorem 3.5 $G*\mathbb{Z}$ is existentially $C^{*}$- residually-$G$, so that by Corollary 3.8 $C^{*}_{r}(G)$ has strict comparison, and is in fact selfless. ∎
3.4. On the particular case of free groups
After initial versions of our paper were released, M. Magee pointed out to us another proof via [46] in the case of non-abelian free groups (avoiding Proposition 3.2). We thank him sincerely for allowing us to include this here.
Theorem 3.10 ([46]).
Let $n>1$ and $a\in\mathbb{F}_{n}$ be a non identity element. Denote by $G=\mathbb{F}_{n}*_{\langle a\rangle}\langle a\rangle\times\mathbb{Z}$, also known as an extension of centralizers of $a$. Then $G$ is existentially $C^{*}$-residually-$\mathbb{F}_{n}$.
Proof.
We consider the maps defined in the proof of Proposition 1.8 of [46], proved using their quantitative Baumslag lemma. Within the proof, note that the maps $f$ constructed are actually identity on the base free group $\mathbb{F}_{n}$ (they are moreover composition of an automorphism fixing $\mathbb{F}_{n}$ and natural retract to $\mathbb{F}_{n}$). Then by following Theorem 1.6 of [46], the desired conclusion is reached. ∎
Note that for the applications described in this paper (via applying Proposition 3.7), we need the degenerate case where $a=1_{\mathbb{F}_{n}}$ in the above (which is for instance handled directly by our Proposition 3.2). While this is not handled directly in the above argument, it is possible to deduce this case via the following observation: take $n=2$, $\mathbb{F}_{2}=\langle x,y\rangle$ and $a=y$, in the above Theorem. Then clearly, $G\cong\mathbb{Z}*(\mathbb{Z}\times\mathbb{Z})$, and let the generator commuting with $y$ be denoted $t$. Then the group generated by $x,y,txt$ is isomorphic to $\mathbb{F}_{3}$ as a subgroup of $G$, and contains the original $\mathbb{F}_{2}$. This can therefore serve to replace $G$ in the conclusion of the Theorem as needed. An identical argument works for any $n>2$ as well.
@@ Appendix A Quantitative version of Wenyuan Yang’s admissible path lemma
The aim of this section is to prove a quantitative version of W. Yang’s admissible path lemma [72, Corollary 3.4], which applies to our setting. For our purposes, we will take a fixed action on a hyperbolic space but vary the loxodromic element and thus the quasi-geodesic constants $\lambda$. So throughout, we track dependence on $\lambda$. The following quantitative version of the Morse lemma is present in [31, Section 5.2]; see [63, Theorem 1.1] for an explicit polynomial.
@@
$g]=\inf_{x\in X}d(gx,x).$
This infimum is in fact achieved at some point $x_{0}\in X$ (see the discussion after Definition 5 in [24]). We let $L_{g}=\cup_{n\in{\mathbb{Z}}}[g^{n}x_{0},g^{n+1}x_{0}]$. Let $\gamma=\cup_{n\in{\mathbb{Z}}}[g^{n}o,g^{n+1}o]$ be the axis of $g$ going through $o.$ Let $\lambda=\frac{d(go,o)}{\tau(g)}$ so that $\gamma$ is a $(\lambda,0)$-quasi-axis; i.e., the length of any sub-segment of $\gamma$ is at most $\lambda$ times the distance between the endpoints. That is, if $m<n$ are integers then
$\sum_{i=m}^{n-1}d(g^{i}o,g^{i+1}o)\leq=\lambda d(g^{n-m}o,o).$
Note that clearly $g]\leq d(go,o)=\tau(g)\lambda.$ Let $E(g)$ denote the maximal elementary (i.e., virtually cyclic) subgroup of $G$ containing $g.$ For sets $Y,Z\subset X$, let $\pi_{Y}(Z)$ denote the projection of $Z$ onto $Y.$ More precisely, $y\in\pi_{Y}(Z)$ if and only if $y\in Y$ and there exists $z\in Z$ such that $d(y,z)=\min\{d(y^{\prime},z):y^{\prime}\in Y\}$. For a set $Y\subset X$ let $\|Y\|$ denote its diameter; i.e., the smallest real number such that $d(y,y^{\prime})\leq\|Y\|$ for all $y,y^{\prime}\in Y.$
The following is due to Delzant.
@@
([24, Proposition 5]) Let $G\curvearrowright X$ be an acylindrical action on a hyperbolic space. Then there exists a constant $D_{0}$ such that for all loxodromic elements (i.e. hyperbolic isometries) $g\in G$ and for all $h\not\in E(g)$,
@@
The following lemma is the quantitative version of [68, Lemma 2.23].
@@
Let $G\curvearrowright X$ be an acylindrical action on a hyperbolic space. Then there is a quadratic $Q_{2}$ such that whenever a loxodromic element $g$ and a point $x\in X$ gives a $(\lambda,0)$-quasi-axis $\gamma=\cup_{n\in{\mathbb{Z}}}[g^{n}x,g^{n+1}x]$ and $h\not\in E(g),$ then
@@
The following theorem, the main theorem of this section, is essentially a reproduction of [72, Corollary 3.4] but where we have tracked all of the constants and dependencies to verify they grow polynomially.
@@
Set $\lambda=\frac{d(go,o)}{\tau(g)}$. Let $\mathcal{L}$} be the collection of $(\lambda,0)$-quasi-geodesics in $X$; in particular, $\mathcal{L}$} includes all geodesics since $\lambda\geq 1$. Let $\mathbb{X}$} be defined by ${\mathbb{X}}=\{\cup_{n=k}^{\ell}[hg^{n}o,hg^{n+1}o]:h\not\in E(g)\}$ and note that $\mathbb{X}$$\subset{\mathcal{L}}$. By Lemma A.1 and the proof of [10, Corollary 3.4], we have that $\mathbb{X}$} is $(\mu,\varepsilon)$-contracting with respect to $\mathcal{L}$} where $\mu,\varepsilon$ are uniformly bounded above by a quadratic $Q_{3}(\lambda)$. That is, whenever $q\in{\mathcal{L}}$ and $p\in{\mathbb{X}}$ such that $d(q,p)\geq Q_{3}(\lambda)$ then $\|\pi_{p}(q)\|<Q_{3}(\lambda)$. (In the notation of [72], $\mu(\lambda^{\prime},c^{\prime}),\varepsilon(\lambda^{\prime},c^{\prime})\leq Q_{3}(\lambda)$ for all $\lambda^{\prime},c^{\prime}\in{\mathbb{R}}$.)
We now obtain that for each $\in{\mathbb{X}}, pp$ is $\sigma$-quasiconvex where $\sigma(U)=\frac{3}{2}Q_{3}(\lambda)+2U$. Indeed, let $q$ be a geodesic in $X$ such that its endpoints $q^{+}$ and $q^{-}$ are in the $U$-neighborhood $N_{U}(p)$ of $p.$ Since $p$ is $(\mu,\varepsilon)$-contracting, either $d(q,p)<\mu$ (and we are done) or $\|\pi_{p}(q)\|<\varepsilon$. In particular, we have $d(q^{+},q^{-})\leq\varepsilon+2U$. Since $q$ is a geodesic, for any $x\in q$ we have $d(q^{+},x)+d(q^{-},x)=d(q^{+},q^{-})$. Therefore without loss of generality $d(q^{+},x)\leq\varepsilon/2+U$, implying that $d(x,\pi_{p}(x))<2U+\frac{3}{2}\varepsilon$. Thus $q\subset N_{\sigma(U)}(p)$.
By Lemma A.3, if $p,p^{\prime}\in{\mathbb{X}}$ are quasi-geodesics corresponding to $h,h^{\prime}$ in different left cosets of $E(g)$, then $p,p^{\prime}$ have bounded projection $T$ where $T=Q_{2}(\lambda)(1+[g])=O(\lambda^{3})$ since $g]\leq d(go,o)=\tau(g)\lambda$ (and the stable length $\tau(g)$ is fixed). We note that in [72], the symbol $\tau$ is used for bounded projection but we have already used $\tau$ for stable length. By [72, Lemma 2.7], $p,p^{\prime}$ also have $\nu$-bounded intersection where $\nu(U)=2U+O(\lambda^{3})$.
@@
$P$ $=[o,h_{1}o]\cup[h_{1}o,h_{1}go]\cup\ldots\cup[h_{1}g^{n_{1}-1}o,h_{1}g^{n_{1}}o]\cup[h_{1}g^{n_{1}}o,h_{1}g^{n_{1}}h_{2}o]\cup\ldots$
$\ldots\cup[h_{1}g^{n_{1}}\cdots h_{k}g^{n_{k}-1}o,h_{1}g^{n_{1}}\cdots h_{m}g^{n_{m}}o];$
set $q_{i}$ to be the geodesic subsegment of $P$ where $h_{i}$ first appears and set $p_{i}$ to be the subpath of $P$ between $q_{i}$ and $q_{i+1}$ (with $p_{m}$ being the subpath of $P$ coming after $q_{m}$). More explicitly, for $i=1,2$ we have that $q_{i}$ and $p_{i}$ are as follows:
@@
We are now ready to analyze the constants in the proof of [72, Proposition 3.3]. We have $c=0$ everywhere.
@@
$B_{\lambda,0}$ $=2\varepsilon(1,0)+2\mu(1,0)+\nu(\mu(1,0)+\sigma(0))+C^{\prime}|g|_{S}$
@@
$\quad\quad\quad\ \ 2C^{\prime}|g|_{S}+3\varepsilon(1,0)+6\mu(1,0),$
@@
Now, if we choose all of the $n_{i}$ so that $|n_{i}|$ are larger than $\lambda D=O(|g|_{S}^{6}),$ we can apply [72, Corollary 3.4] to obtain that our path $P$ is a $(\Lambda,0)$-quasi-geodesic. In particular, $P$ is nontrivial and so it must be that $h_{1}g^{n_{1}}h_{2}g^{n_{2}}\cdots h_{m}g^{n_{m}}\neq e$. (Note that in the case $n_{m}=0$, we can still apply [72, Corollary 3.4].) This finishes the proof. ∎